# Representing in Low Rank I: conjugacy, topological and homological aspects

### v1

@@ 3.2. Division algebra factors in case of property $\operatorname{(M_{exc})}$
Thus (3.3) can be reformulated as saying that $FG$ is covered by $\{Ge\mid e\in\operatorname{PCI}(FG)\}$. To obtain the restrictions on the simple components of $FG$, we will describe which finite groups can be isomorphic to $Ge$ with $e\in\operatorname{PCI}(FG)$ when $FG$ has $\operatorname{(M_{exc})}$. In [CdRZ, Proposition 4.1] such a study was essentially already done. The proof in loc. cit. however contains some flaws, which we correct. Note that if $FG$ has $\operatorname{(M_{exc})}$, then each $Ge$ is a subgroup of either a division algebra or of an exceptional matrix algebra. We will call finite subgroups of such simple algebras low rank spanning groups.
@@
    If ${\mathbb{Q}}G$ has $\operatorname{(M_{exc})}$, then the only possible exceptional matrix algebras [^6] Note that by [EKVG, Theorem 3.5] the only exceptional matrix algebra not appearing is $\operatorname{M}_{2}(\left(\frac{-2,-5}{{\mathbb{Q}}}\right))$. are
@@
In the proof we will obtain that for $C_{3}\rtimes C_{2^{n}}$, with $C_{2^{n}}$ acting by inversion, the rational group algebra has all non-division simple components of the form $\operatorname{M}_{2}({\mathbb{Q}})$ and $\operatorname{M}_{2}({\mathbb{Q}}(i))$. Furthermore, it has a division component of the form $\left(\frac{\zeta_{2^{n-1}},-3}{{\mathbb{Q}}(\zeta_{2^{n-1}})}\right)$. Hence it gives an example of a group that has all matrix components of the form $\operatorname{M}_{2}({\mathbb{Q}}(\sqrt{-d}))$ with $d\in{\mathbb{N}}$ but which does not satisfy the stronger property considered in [FreebyFree], where they require that ${\mathbb{Q}}G$ has no exceptional division components.
One interesting class of groups for which ${\mathbb{Q}}G$ has no exceptional division components is the class of cut groups, see [BJJKT, Proposition 6.12]. Recall that a group is called cut if $\mathcal{Z}({\mathcal{U}}({\mathbb{Z}}G))$ is finite.
@@ 3.2.1. Background on strong Shoda pairs and simple components
We need to recall some methods to construct primitive central idempotents of ${\mathbb{Q}}G$. These methods were introduced by Olivieri–del Río–Simón [OlRiSi], see [EricAngel1, Chapter 3] for a good introduction.
@@
The following is a combination of [EricAngel1, Proposition 3.4.1, Theorems 3.4.2 & 3.5.5 and Problem 3.5.1] and [JM, Lemma 3.4].
Theorem 3.7 ([OlRiSi]).
@@ 3.2.2. Proof of the main result
Using Lemma˜3.9, the proof of Theorem˜3.4 reduces to a study of the subgroups of division algebras classified by Amitsur [Amitsur] and the groups spanning an exceptional matrix algebra classified in [EKVG].
@@
We start with $\operatorname{PCI}_{2}$. The finite subgroups $\mathcal{G}$ of $\operatorname{GL}_{2}({\mathbb{Q}}(\sqrt{-d}))$ or $\operatorname{GL}_{2}(\left(\frac{-a,-b}{{\mathbb{Q}}}\right))$ with $a,b,d\in{\mathbb{N}}$ with the property that $\operatorname{Span}_{{\mathbb{Q}}}\{\mathcal{G}\}$ is the respective $\operatorname{M}_{2}(\cdot)$ have been classified [^7] In [EKVG, Table 2] a group was missing, see [BJJKT, Appendix A] for a complete list. in [EKVG, Theorem 3.7]. This classification consists of $55$ groups and in particular the groups $Ge$ for $e\in\operatorname{PCI}_{2}$ are among them. The simple components of ${\mathbb{Q}}\mathcal{G}$ may be computed using e.g. the Wedderga package in GAP, see Table˜B in Appendix˜A for the result. This moreover shows that for groups $Ge$ with $e\in\operatorname{PCI}_{2}$, the property $\operatorname{(M_{exc})}$ is equivalent to the weaker property that each non-division component has reduced degree $2$. A case-by-case verification also shows that each of these groups $Ge$ contains an abelian normal subgroup of index a divisor of $4$. Furthermore, as can be seen in the table, the only $1\times 1$ components appearing are
@@
Next we consider the case $e\in\operatorname{PCI}_{1}$. A similar reasoning applies. Indeed, the finite subgroups (such as $Ge$ for $e\in\operatorname{PCI}_{1}$) of rational division algebras have been classified by Amitsur in [Amitsur]. We will use the rephrasing from [ShiWeh, Theorem $2.1.4$] which asserts that they are:
@@
This is a metabelian group and hence its strong Shoda pairs are described by Theorem˜3.8, which we apply now. Note that the derived subgroup $Dic_{4n}^{\prime}$ is given by $\langle a^{2}\rangle$, and $\langle a\rangle$ is the maximal abelian subgroup of $Dic_{4n}$ containing it. For any SSP $(H,K)$, one has that $\langle a\rangle\subseteq H$. In other words, $H=\langle a\rangle$ or $H=Dic_{4n}$. If $H=Dic_{4n}$ then the simple component associated to $(H,K)$ is a field, by [JespersSun, Lemma 2.4]. Via Theorem˜3.8, it is a direct verification that for $d\mid 2n$, the tuple $(\langle a\rangle,\langle a^{d}\rangle)$ is a SSP if and only if $d\neq 1,2$. Note that $K$ is normal in $Dic_{4n}$. Hence the associated primitive central idempotent is $\epsilon(\langle a\rangle,\langle a^{d}\rangle)$. Now, in [EricAngel1, Example $3.5.7$], it is noted that (for $n$ not necessarily even)
@@
Before we consider case (b)(iii) from above, we will discuss (a), the Z-groups (and in particular $Dic_{4.3}$). They have also been classified, see [ShiWeh, Theorem $2.1.5$]. They are the following:
@@
It is clear that the cyclic groups have $\operatorname{(M_{exc})}$ since ${\mathbb{Q}}C_{n}$ is abelian. Moreover, by the theorem of Perlis–Walker [EricAngel1, Theorem 3.3.6], $\mathcal{C}({\mathbb{Q}}C_{n})=\{{\mathbb{Q}}(\zeta_{d})\mid d\text{ divides }n\}$.
@@
where $k_{j}:=b_{j}-\alpha_{j}>0$ and the action is non-trivial and faithful. Denote by $x$ and $y$ the respective generators of the factors of $B$, i.e. $B=\langle x\rangle\rtimes\langle y\rangle$. By Lemma˜3.9, the group $B$ also has $\operatorname{(M_{exc})}$. Note that $C_{p^{a}}=\langle x\rangle$ is a maximal abelian subgroup of $B$ containing $B^{\prime}$. Now using Theorem˜3.8, it is a direct verification that $(H,K)=(\langle x\rangle,1)$ is a SSP of $B$. Moreover, ${\mathbb{Q}}Be(G,\langle x\rangle,1)\cong{\mathbb{Q}}(\zeta_{o(x)})*\langle y\rangle$ for some explicit crossing (see [EricAngel1, Remark 3.5.6]) that would show that the component is not a division algebra. Hence, as $B$ has $\operatorname{(wM_{exc})}$, it is an exceptional matrix algebra or $\operatorname{M}_{4}({\mathbb{Q}})$ and thus $\dim_{{\mathbb{Q}}}{\mathbb{Q}}Be(G,\langle x\rangle,1)\mid 16$. Now using [JM, Lemma 3.4], we compute that
@@
As $m$ and $o_{m}(2)$ are odd, all the components are division algebras – a conclusion that also directly would have followed from [SehNilpotent]. In conclusion [^11] This conclusion only requires that every non-division component is of the form $\operatorname{M}_{2}(D)$ with $\mathcal{Z}(D):{\mathbb{Q}}]\leq 2$, not the full strength of $\operatorname{(M_{exc})}$.,
@@ 4.1. Groups of virtual cohomological dimension 4
Recall that in [FreebyFree, Proposition 3.3] the finite-dimensional simple $F$-algebras AA, over a number field $F$, such that $\operatorname{vcd}(\operatorname{SL}_{1}(\mathcal{O}))\leq 2$ for some (and hence each) order $\mathcal{O}$ in $A$ have been classified. We extend this result to virtual cohomological dimension $4$. See Remark˜4.4 below for $\operatorname{vcd}(\operatorname{SL}_{1}(\mathcal{O}))=3$.
@@
Let $D$ a division ring of degree $d$, for an integer $n\geq 1$, and let $A=M_{n}(D)$, and $F=\mathcal{Z}(D)$. Let $\mathcal{O}$ be an order in $A$. We make use of the following formula, as stated in [FreebyFree, Eq. (1)].
@@
where $s$ is the number of pairs of non-real complex embeddings of $F$, $r_{1}$ is the number of real embeddings of $F$ at which $A$ is ramified, and $r_{2}$ the number of real embeddings of $F$ at which $A$ is not ramified. We may assume that $nd>1$, since when $nd=1$, AA is a field, which implies that $\operatorname{vcd}(\operatorname{SL}_{1}(\mathcal{O}))=0$ by [FreebyFree, Proposition 3.3]. Note that for any choice of $nd\geq 2$, the first two terms of eq.˜4.1 are non-negative. Additionally, when $d$ is odd, it is well-known that $r_{1}=0$.
@@
Concerning $\operatorname{vcd}(\operatorname{SL}_{1}(\mathcal{O}))\leq 2$, [FreebyFree, Proposition 3.3] shows that
@@
From [Koo, Page 147], it follows that for any $g\in Ge$ of prime power order $p^{k}$, the $p^{k}$th cyclotomic polynomial $\Phi_{p^{k}}$ equals $\mu_{{\mathbb{Q}},g}$. In particular, ${\mathbb{Q}}(\zeta_{p^{k}})\cong\frac{{\mathbb{Q}}[X]}{(\mu_{{\mathbb{Q}},g})}$. The latter also holds over $F$, when $p^{k}>4$. Indeed, $\mu_{F,g}$ is given as the unique monic polynomial generating the ideal $I_{g}:=\{P\in F[X]\mid P(g)=0\}$, and the minimal polynomial of $\zeta_{p^{k}}$ over $F$ is given by the unique monic polynomial generating the ideal $I_{\zeta_{p^{k}}}:=\{Q\in F[X]\mid Q(\zeta_{p^{k}})=0\}$. We claim that $I_{g}=I_{\zeta_{p^{k}}}$. Indeed, by [Koo, Page 146], there is some matrix $B\in\operatorname{GL}_{2}({\mathbb{C}})$ such that $BgB^{-1}={\rm diag}(\lambda_{1},\lambda_{2})$, with each $\lambda_{i}$ a $p^{k}$th root of unity, amongst which at least one primitive (otherwise $BgB^{-1}$ and hence $g$ would have order strictly smaller than $p^{k}$). Without loss of generality, assume $\lambda_{1}=\zeta_{p^{k}}$. Remark that since conjugation by an invertible matrix induces an algebra automorphism of $\operatorname{M}_{2}({\mathbb{C}})$, it follows that $P\in I_{g}$ if and only if $P(BgB^{-1})=0$. In particular, $P\in I_{g}$ if and only if $P(\zeta_{p^{k}})=0=P(\lambda_{2})$. We conclude that $I_{g}\subseteq I_{\zeta_{p^{k}}}$. But since $\deg(\mu_{F,g})=2$, and $p^{k}>4$ (meaning that $\zeta_{p^{k}}\not\in F$), it follows that $I_{g}$ is a maximal ideal of $F[X]$, and hence $I_{g}=I_{\zeta_{p^{k}}}$. In particular, $F(\zeta_{p^{k}})\cong\frac{F[X]}{(\mu_{F,g})}$.
@@
Let now $\pi(Ge)\subseteq\{2,3\}$. We bound the exponent of $Ge$. Let $g\in Ge$, say of order $2^{i}3^{j}$ for some non-negative integers $i$, $j$. If $i=0$ or $j=0$, then $o(g)=p^{n}$ with $p\in\{2,3\}$. Since $\Phi_{p^{n}}$ divides $\mu_{{\mathbb{Q}},g}$ for each $p\in\{2,3\}$ and $\deg(\mu_{{\mathbb{Q}},g})\leq 6$, it follows that $o(g)\in\{p,p^{2}\}$. If $i\neq 0\neq j$, then considering a realisation of $g$ as an element in $\operatorname{GL}_{6}({\mathbb{Q}})$, [Koo, Page 147] implies that there exist positive integers $m_{1},\ldots,m_{r}$ such that $o(g)=\mathrm{lcm}\{m_{1},\ldots,m_{r}\}$, $\Phi_{m_{i}}$ divides $\mu_{{\mathbb{Q}},g}$, and $6=\sum_{i=1}^{r}d_{i}\phi(m_{i})$, for some $d_{i}\geq 1$. In particular each $\phi(m_{i})\leq 6$. Since $\pi(Ge)\subseteq\{2,3\}$, from a case-by-case analysis it follows that $m_{i}\in\{1,2,3,4,6,9,18\}$. In particular, $\exp(G)\mid 36$.

### v2

@@ 3.2. Division algebra factors in case of property $\operatorname{(M_{exc})}$
Thus (3.3) can be reformulated as saying that $FG$ is covered by $\{Ge\mid e\in\operatorname{PCI}(FG)\}$. To obtain the restrictions on the simple components of $FG$, we will describe which finite groups can be isomorphic to $Ge$ with $e\in\operatorname{PCI}(FG)$ when $FG$ has $\operatorname{(M_{exc})}$. In [13, Proposition 4.1] such a study was essentially already done. The proof in loc. cit. however contains some flaws, which we correct. Note that if $FG$ has $\operatorname{(M_{exc})}$, then each $Ge$ is a subgroup of either a division algebra or of an exceptional matrix algebra. We will call finite subgroups of such simple algebras low rank spanning groups.
@@
    If ${\mathbb{Q}}G$ has $\operatorname{(M_{exc})}$, then the only possible exceptional matrix algebras [^6] Note that by [23, Theorem 3.5] the only exceptional matrix algebra not appearing is $\operatorname{M}_{2}(\left(\frac{-2,-5}{{\mathbb{Q}}}\right))$. are
@@
In the proof we will obtain that for $C_{3}\rtimes C_{2^{n}}$, with $C_{2^{n}}$ acting by inversion, the rational group algebra has all non-division simple components of the form $\operatorname{M}_{2}({\mathbb{Q}})$ and $\operatorname{M}_{2}({\mathbb{Q}}(i))$. Furthermore, it has a division component of the form $\left(\frac{\zeta_{2^{n-1}},-3}{{\mathbb{Q}}(\zeta_{2^{n-1}})}\right)$. Hence it gives an example of a group that has all matrix components of the form $\operatorname{M}_{2}({\mathbb{Q}}(\sqrt{-d}))$ with $d\in{\mathbb{N}}$ but which does not satisfy the stronger property considered in [50], where they require that ${\mathbb{Q}}G$ has no exceptional division components.
One interesting class of groups for which ${\mathbb{Q}}G$ has no exceptional division components is the class of cut groups, see [4, Proposition 6.12]. Recall that a group is called cut if $\mathcal{Z}({\mathcal{U}}({\mathbb{Z}}G))$ is finite.
@@ 3.2.1. Background on strong Shoda pairs and simple components
We need to recall some methods to construct primitive central idempotents of ${\mathbb{Q}}G$. These methods were introduced by Olivieri–del Río–Simón [67], see [47, Chapter 3] for a good introduction.
@@
The following is a combination of [47, Proposition 3.4.1, Theorems 3.4.2 & 3.5.5 and Problem 3.5.1] and [62, Lemma 3.4].
Theorem 3.7 ([67]).
@@ 3.2.2. Proof of the main result
Using Lemma˜3.9, the proof of Theorem˜3.4 reduces to a study of the subgroups of division algebras classified by Amitsur [2] and the groups spanning an exceptional matrix algebra classified in [23].
@@
We start with $\operatorname{PCI}_{2}$. The finite subgroups $\mathcal{G}$ of $\operatorname{GL}_{2}({\mathbb{Q}}(\sqrt{-d}))$ or $\operatorname{GL}_{2}(\left(\frac{-a,-b}{{\mathbb{Q}}}\right))$ with $a,b,d\in{\mathbb{N}}$ with the property that $\operatorname{Span}_{{\mathbb{Q}}}\{\mathcal{G}\}$ is the respective $\operatorname{M}_{2}(\cdot)$ have been classified [^7] In [23, Table 2] a group was missing, see [4, Appendix A] for a complete list. in [23, Theorem 3.7]. This classification consists of $55$ groups and in particular the groups $Ge$ for $e\in\operatorname{PCI}_{2}$ are among them. The simple components of ${\mathbb{Q}}\mathcal{G}$ may be computed using e.g. the Wedderga package in GAP, see Table˜B in Appendix˜A for the result. This moreover shows that for groups $Ge$ with $e\in\operatorname{PCI}_{2}$, the property $\operatorname{(M_{exc})}$ is equivalent to the weaker property that each non-division component has reduced degree $2$. A case-by-case verification also shows that each of these groups $Ge$ contains an abelian normal subgroup of index a divisor of $4$. Furthermore, as can be seen in the table, the only $1\times 1$ components appearing are
@@
Next we consider the case $e\in\operatorname{PCI}_{1}$. A similar reasoning applies. Indeed, the finite subgroups (such as $Ge$ for $e\in\operatorname{PCI}_{1}$) of rational division algebras have been classified by Amitsur in [2]. We will use the rephrasing from [83, Theorem $2.1.4$] which asserts that they are:
@@
This is a metabelian group and hence its strong Shoda pairs are described by Theorem˜3.8, which we apply now. Note that the derived subgroup $Dic_{4n}^{\prime}$ is given by $\langle a^{2}\rangle$, and $\langle a\rangle$ is the maximal abelian subgroup of $Dic_{4n}$ containing it. For any SSP $(H,K)$, one has that $\langle a\rangle\subseteq H$. In other words, $H=\langle a\rangle$ or $H=Dic_{4n}$. If $H=Dic_{4n}$ then the simple component associated to $(H,K)$ is a field, by [45, Lemma 2.4]. Via Theorem˜3.8, it is a direct verification that for $d\mid 2n$, the tuple $(\langle a\rangle,\langle a^{d}\rangle)$ is a SSP if and only if $d\neq 1,2$. Note that $K$ is normal in $Dic_{4n}$. Hence the associated primitive central idempotent is $\epsilon(\langle a\rangle,\langle a^{d}\rangle)$. Now, in [47, Example $3.5.7$], it is noted that (for $n$ not necessarily even)
@@
Before we consider case (b)(iii) from above, we will discuss (a), the Z-groups (and in particular $Dic_{4.3}$). They have also been classified, see [83, Theorem $2.1.5$]. They are the following:
@@
It is clear that the cyclic groups have $\operatorname{(M_{exc})}$ since ${\mathbb{Q}}C_{n}$ is abelian. Moreover, by the theorem of Perlis–Walker [47, Theorem 3.3.6], $\mathcal{C}({\mathbb{Q}}C_{n})=\{{\mathbb{Q}}(\zeta_{d})\mid d\text{ divides }n\}$.
@@
where $k_{j}:=b_{j}-\alpha_{j}>0$ and the action is non-trivial and faithful. Denote by $x$ and $y$ the respective generators of the factors of $B$, i.e. $B=\langle x\rangle\rtimes\langle y\rangle$. By Lemma˜3.9, the group $B$ also has $\operatorname{(M_{exc})}$. Note that $C_{p^{a}}=\langle x\rangle$ is a maximal abelian subgroup of $B$ containing $B^{\prime}$. Now using Theorem˜3.8, it is a direct verification that $(H,K)=(\langle x\rangle,1)$ is a SSP of $B$. Moreover, ${\mathbb{Q}}Be(G,\langle x\rangle,1)\cong{\mathbb{Q}}(\zeta_{o(x)})*\langle y\rangle$ for some explicit crossing (see [47, Remark 3.5.6]) that would show that the component is not a division algebra. Hence, as $B$ has $\operatorname{(wM_{exc})}$, it is an exceptional matrix algebra or $\operatorname{M}_{4}({\mathbb{Q}})$ and thus $\dim_{{\mathbb{Q}}}{\mathbb{Q}}Be(G,\langle x\rangle,1)\mid 16$. Now using [62, Lemma 3.4], we compute that
@@
As $m$ and $o_{m}(2)$ are odd, all the components are division algebras – a conclusion that also directly would have followed from [79]. In conclusion [^11] This conclusion only requires that every non-division component is of the form $\operatorname{M}_{2}(D)$ with $\mathcal{Z}(D):{\mathbb{Q}}]\leq 2$, not the full strength of $\operatorname{(M_{exc})}$.,
@@ 4.1. Groups of virtual cohomological dimension 4
Recall that in [50, Proposition 3.3] the finite-dimensional simple $F$-algebras AA, over a number field $F$, such that $\operatorname{vcd}(\operatorname{SL}_{1}(\mathcal{O}))\leq 2$ for some (and hence each) order $\mathcal{O}$ in $A$ have been classified. We extend this result to virtual cohomological dimension $4$. See Remark˜4.4 below for $\operatorname{vcd}(\operatorname{SL}_{1}(\mathcal{O}))=3$.
@@
Let $D$ a division ring of degree $d$, for an integer $n\geq 1$, and let $A=M_{n}(D)$, and $F=\mathcal{Z}(D)$. Let $\mathcal{O}$ be an order in $A$. We make use of the following formula, as stated in [50, Eq. (1)].
@@
where $s$ is the number of pairs of non-real complex embeddings of $F$, $r_{1}$ is the number of real embeddings of $F$ at which $A$ is ramified, and $r_{2}$ the number of real embeddings of $F$ at which $A$ is not ramified. We may assume that $nd>1$, since when $nd=1$, AA is a field, which implies that $\operatorname{vcd}(\operatorname{SL}_{1}(\mathcal{O}))=0$ by [50, Proposition 3.3]. Note that for any choice of $nd\geq 2$, the first two terms of eq.˜4.1 are non-negative. Additionally, when $d$ is odd, it is well-known that $r_{1}=0$.
@@
Concerning $\operatorname{vcd}(\operatorname{SL}_{1}(\mathcal{O}))\leq 2$, [50, Proposition 3.3] shows that
@@
From [57, Page 147], it follows that for any $g\in Ge$ of prime power order $p^{k}$, the $p^{k}$th cyclotomic polynomial $\Phi_{p^{k}}$ equals $\mu_{{\mathbb{Q}},g}$. In particular, ${\mathbb{Q}}(\zeta_{p^{k}})\cong\frac{{\mathbb{Q}}[X]}{(\mu_{{\mathbb{Q}},g})}$. The latter also holds over $F$, when $p^{k}>4$. Indeed, $\mu_{F,g}$ is given as the unique monic polynomial generating the ideal $I_{g}:=\{P\in F[X]\mid P(g)=0\}$, and the minimal polynomial of $\zeta_{p^{k}}$ over $F$ is given by the unique monic polynomial generating the ideal $I_{\zeta_{p^{k}}}:=\{Q\in F[X]\mid Q(\zeta_{p^{k}})=0\}$. We claim that $I_{g}=I_{\zeta_{p^{k}}}$. Indeed, by [57, Page 146], there is some matrix $B\in\operatorname{GL}_{2}({\mathbb{C}})$ such that $BgB^{-1}={\rm diag}(\lambda_{1},\lambda_{2})$, with each $\lambda_{i}$ a $p^{k}$th root of unity, amongst which at least one primitive (otherwise $BgB^{-1}$ and hence $g$ would have order strictly smaller than $p^{k}$). Without loss of generality, assume $\lambda_{1}=\zeta_{p^{k}}$. Remark that since conjugation by an invertible matrix induces an algebra automorphism of $\operatorname{M}_{2}({\mathbb{C}})$, it follows that $P\in I_{g}$ if and only if $P(BgB^{-1})=0$. In particular, $P\in I_{g}$ if and only if $P(\zeta_{p^{k}})=0=P(\lambda_{2})$. We conclude that $I_{g}\subseteq I_{\zeta_{p^{k}}}$. But since $\deg(\mu_{F,g})=2$, and $p^{k}>4$ (meaning that $\zeta_{p^{k}}\not\in F$), it follows that $I_{g}$ is a maximal ideal of $F[X]$, and hence $I_{g}=I_{\zeta_{p^{k}}}$. In particular, $F(\zeta_{p^{k}})\cong\frac{F[X]}{(\mu_{F,g})}$.
@@
Let now $\pi(Ge)\subseteq\{2,3\}$. We bound the exponent of $Ge$. Let $g\in Ge$, say of order $2^{i}3^{j}$ for some non-negative integers $i$, $j$. If $i=0$ or $j=0$, then $o(g)=p^{n}$ with $p\in\{2,3\}$. Since $\Phi_{p^{n}}$ divides $\mu_{{\mathbb{Q}},g}$ for each $p\in\{2,3\}$ and $\deg(\mu_{{\mathbb{Q}},g})\leq 6$, it follows that $o(g)\in\{p,p^{2}\}$. If $i\neq 0\neq j$, then considering a realisation of $g$ as an element in $\operatorname{GL}_{6}({\mathbb{Q}})$, [57, Page 147] implies that there exist positive integers $m_{1},\ldots,m_{r}$ such that $o(g)=\mathrm{lcm}\{m_{1},\ldots,m_{r}\}$, $\Phi_{m_{i}}$ divides $\mu_{{\mathbb{Q}},g}$, and $6=\sum_{i=1}^{r}d_{i}\phi(m_{i})$, for some $d_{i}\geq 1$. In particular each $\phi(m_{i})\leq 6$. Since $\pi(Ge)\subseteq\{2,3\}$, from a case-by-case analysis it follows that $m_{i}\in\{1,2,3,4,6,9,18\}$. In particular, $\exp(G)\mid 36$.